OpenStudy (toxicsugar22):

9y^2-8x^2+72y+16x+64=0 give the exact coordinates of any four points on the hyperbola

OpenStudy (toxicsugar22):

hi can u help me

OpenStudy (anonymous):

let y be 0 8x^2 - 16x - 64 = 0 x^2 - 2x - 8 = 0 x = 4 or -2 therefore 2 points are (-2,0) and (4,0) now let x be 0 9y^2 + 72y + 64 = 0 y = -1.01 or -6.98 therefore the next 2 points are (0,-1.01) and (0, -6.98)

OpenStudy (toxicsugar22):

i dont think the (0,-1.01) and (0, -6.98) are correct

OpenStudy (toxicsugar22):

i think we should let y be something and solve for x